3x-7y=2
-5x+4y=12

Answer by checkley79(3341)

3x-7y=2 multiply by 5
-5x+4y=12 multiply by 3 & add.
#########################
15x-35y=10
-15x+12y=36
-----------------
-23y=46
y=46/-23
y=-2 ans.
3x-7*-2=2
3x+14=2
3x=2-14
3x=-12
x=-12/3
x=-4 ans.
Proof:
-5(-4)+4(-2)=12
20-8=12
12=12